Carney college is expanding from 850 to 1050 students. To avoid any rise in the student to faculty ratio, the faculty of 69 professors must also iincrease. How many new faculty positions should be created?
Answer by edjones(8007)

x/69=1050/850
69x/69=1050*69/850
x=1050*69/850
x=85.2353
since we can't have a portion of a professor we round to 85.
Ed